OpenStudy (jocelynstalnaker):

Anyone willing to help me? Write the equation of the line that is parallel to the line 4x - 3y = -12 and passes through the point (-3, 4). A. y = four thirdsx + 8 B. y = four thirdsx + 3 C. y = -three fourthsx + 8 D. y = -three fourthsx + 3

OpenStudy (jocelynstalnaker):

@tkhunny

OpenStudy (jocelynstalnaker):

@ikram002p

OpenStudy (jocelynstalnaker):

@nikato

OpenStudy (nikato):

you should know that if two lines are parallel, that means they have the same slope. so lets find the slope of 4x-3y=-12 do you know how?

OpenStudy (jocelynstalnaker):

Yes, well tell me if i'm right. 3y=-12+4x

OpenStudy (jocelynstalnaker):

Then divide 3

OpenStudy (jocelynstalnaker):

y=-4+

OpenStudy (jocelynstalnaker):

y=-4+4/3x?

OpenStudy (nikato):

it shouldve been 3y=12+4x not negative 12 but it wouldnt really affect you answer in this question

OpenStudy (jocelynstalnaker):

Oh ok. so y=4+4/3x

OpenStudy (nikato):

so your equation would be y=4+ 4/3x so whats the slope?

OpenStudy (jocelynstalnaker):

4/3? or 4?

OpenStudy (jocelynstalnaker):

4/3 right?

OpenStudy (nikato):

yea, 4/3 is right remember slope-intercept form which is what you put it in